A block of mass m = 6.20 kg is released from rest
Answer #1

a) We use the conservation of energy

For B:

E_A = E_B

mgh_A = mgh_B+mv_B^2/2

v_B = \sqrt{2g(h_A-h_B)}= \sqrt{2(9.81)(6.1-3.2)}=7.543 m/sv_B = \sqrt{2g(h_A-h_B)}= 7.543 m/s

For C:

E_A = E_C

mgh_A = mgh_C+mv_C^2/2

v_C = \sqrt{2g(h_A-h_C)}= \sqrt{2(9.81)(6.1-2)}=8.969 m/s

b) The work done by gravitational force as it moves from point A to C is:

W_{gravity} = -\Delta U = -(mgh_C-mgh_A) = mg(h_A-h_C)

W_{gravity} = =(6.2)(9.81)(6.1-2)= 249.3702 J